OpenStudy (anonymous):

Determine whether each of the following integrals is convergent or divergent. Evaluate those that are convergent. Intergral 2xe^-x^2 dx from x = infinite to x = 0

OpenStudy (anonymous):

please clarify \[\int_{0}^{\infty} 2xe^{-x^2}\] is this the integral? if so then \[-e^{-x^2} |_{0}^{\infty}\] it is an improper integral you have to write it as a limit, but i will skip that, so \[-e^{-\infty^2}-(-e^{-0^2})=-0+1=1 \]

OpenStudy (anonymous):

I was soooooooooooooooooooooooooooooooo close

OpenStudy (anonymous):

hahahahha, I've had the feeling. Glad to be of help.

OpenStudy (anonymous):

can I give you one more to check

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Well the first one is convergent correct?

OpenStudy (anonymous):

And the next question is intergral e^x from x = 100 to x = infinite

OpenStudy (anonymous):

The first one is convergent, that is correct.

OpenStudy (anonymous):

\[\int_{100}^{\infty} e^x = e^x|_{100}^{\infty}=e^{\infty}-e^{100}=\infty \] so it diverges.

OpenStudy (anonymous):

I'm getting good at this

OpenStudy (anonymous):

That's good news for mathematicians.